A polytropic process of air from 1034 kPa, 149 oC and 0.028 m3 occurs to 138 kPa in accordance with pV1.3 = C. Determine (a) t2 in oC and V2, (b) U and H, (c) Wnf, (d) the heat from the formula using polytropic specific heat.


1
Expert's answer
2021-12-03T12:31:04-0500

a)

"T_2=T_1(\frac{p_2}{p_1})^{1-\frac 1{\gamma}}=265~K=-8°C,"

b)

"A=\frac{p_1V_1}{\gamma-1}(1-(\frac{p_2}{p_1})^{1-\frac 1{\gamma}})=35.8~kPa."